OpenStudy (anonymous):

how do you solve log base 4, if 8 = x

OpenStudy (anonymous):

\[\log_4(8)\] or \[\log_4(x)=8\]?

OpenStudy (anonymous):

first one use \[\log_4(8)=\frac{\log(8)}{\log(4)}\] and a calculator. second one is \[x=4^8\]

OpenStudy (anonymous):

okat thanks!